I am having problems solving this problem: f(x)=1/x, find f(2+h)-f(2)/h
Answer by sdmmadam@yahoo.com(530)

f(x)=1/x, find f(2+h)-f(2)/h
f(2+h)-f(2)/h
=1/(2+h)-[(1/2h)]
[using f(x) = 1/x and f(2)/h means f(2)X(1/h) = (1/2)X(1/h)=1/(2h)]
= [2h-(2+h)]/[(2+h)(2h)]
=[2h-2-h]/2h(2+h)
=(2h-h-2)/2h(2+h)
=(h-2)/2h(2+h)